I wanted to know if y= 6/x + 3 is a function or not??
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)

Is what a function? y=6/x+3 or y=6/(x+3) ?

Answer by ikleyn(52802)

.

Yes, it is.